Consider the economy of Orchard that produces two types of products, apples and oranges. Use the data below to answer the following questions.

                           Apples                    Oranges
A                                 40                              0
B                                 30                              20
C                                 20                              35
D                                 10                              45
E                                  0                                 50

a) Draw the production possibility curve for Orchard.

b) What is the opportunity cost of producing 10 more bushels of apples starting at point B? How does that compare with the cost of 10 more bushels of apples at point D?

c) What would you say about Orchard if it is producing 20 bushels of apples and 30 bushels of oranges?
Answer #1

By plotting the given points ,we get the production possibility curve for Orchard as shown below:

(a) 29 6 Orchards 0 10 20 30 o Apples

(b) The opportunity cost of producing 10 more bushels of apples starting at point B =(20-0)= 20 oranges.

The opportunity cost of producing 10 more bushels of apples starting at point D = (45-35)=10 oranges.

This implies moving from point D to point B , opportunity cost of producing more apples increases.

(c) If Orchard is producing 20 bushels of apples and 30 bushels of oranges ,then this would lie inside the PPF,this means that it is inefficient to produce this combination because with the given inputs, it can produce more of oranges .